OpenStudy (lgbasallote):

what does "obtain a family of solutions" mean in differential equations? it's the instruction in our homework and i dont know what that means lol..is it the particular solution and general solution?

OpenStudy (anonymous):

it means the answer will have a + constant and therefore isn't just one function but a group of functions known as a family. ps can you do this for me? using stokes theorem find the integral of F(x,y,z)=yzi+2xzj+e^xy k where C is the circle x^2+y^2=16 and z=5.

OpenStudy (lgbasallote):

i dont know what that means o.O

OpenStudy (anonymous):

that ok, neither do i.

OpenStudy (anonymous):

hah lgbas? you want to know?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

a family of function happens when you integrate, you know how to integrate correcT?

OpenStudy (lgbasallote):

yeah

OpenStudy (anonymous):

alright so say you integrate something super easy like dx

OpenStudy (anonymous):

\[\int{}{}dx\]=x+c correct?

OpenStudy (lgbasallote):

yes..so x+ c is the family?

OpenStudy (anonymous):

yes because c can be anything say the otherside was dy and you got y=x+c ... this is a family solution when you draw the family c can be 1,2,3,4,5, basically anything; you get something like this